OpenStudy (anonymous):

factor the expression 64a^2-b^2

OpenStudy (scorcher219396):

Its a difference of two squares, something you should recognize and always be looking for when factoring (8a-b)(8a-b) (8a-b)^2

OpenStudy (anonymous):

\[ 64a^2= 8^2a^2 = (8a)^2 \]So we have \[ (8a)^2-b^2 \]Difference of squares.

OpenStudy (anonymous):

Difference of squares says: \[ \color{red}c^2-\color{blue}d^2 = (\color{red}c-\color{blue}d)(\color{red}c+\color{blue}d) \]But in our case we have \(c=8a\) and \(d=b\).\[ (\color{red}{8a})^2-\color{blue}b^2 = (\color{red}{8a}-\color{blue}b)(\color{red}{8a}+\color{blue}b) \]